OpenStudy (anonymous):

Solve by substitution. y=16-5x y=3x

OpenStudy (solomonzelman):

substitute 3y for x into the first equation.

OpenStudy (rina.r):

and will be the 5x for y

OpenStudy (solomonzelman):

y=16-5x y=3x y=16-5x -> y=16-5(3y) -> y=-16-15y y=-16-15y y+15y=-16-15y+15y 16y=-16 y=-1 solve for x (-1)=3x x=-1/3 y=1 x=-1/3

OpenStudy (anonymous):

thank you so much that helped a lot. it says i need an ordered pair. would it just be (-1,3)??

OpenStudy (charlotte123):

@SolomonZelman YOU'RE GOOD! GREAT JOB! :D

OpenStudy (unklerhaukus):

there are a few typos, so be careful , but you final answer , the order pair is the coordinate (x,y)

OpenStudy (anonymous):

soo (-1,1)?

OpenStudy (unklerhaukus):

y=16-5x y=3x sub Eqn 2 into Eqn 1 3x=16-5x = x=

OpenStudy (anonymous):

okay so i'm confused with the ordered pair now

OpenStudy (solomonzelman):

(x,y) if x=-1/3 and y=-3 what is the ordered pair?

OpenStudy (anonymous):

is it (-1/3,-3) could it be a fraction?

OpenStudy (solomonzelman):

i don't know why not.

OpenStudy (shamil98):

y = 16 - 5x y= 3x 3x = 16 - 5x 8x = 16 x = 2 y = 3(2) y = 6. @SolomonZelman you did this wrong.

OpenStudy (anonymous):

Your answer should actually be the ordered pair (2,6). That substitution isn't correct.. \[y=16-5x\]\[y=3x\] \[3x=16-5x\]\[8x=16\]\[x=2\] So \[y=6\] Solution: (2,6)

OpenStudy (solomonzelman):

Sorry guys

OpenStudy (anonymous):

Be really careful with those substitutions to make sure you are making correct mathematical statements!

OpenStudy (shamil98):

\[\huge \frak That~brain~fart.\]

OpenStudy (charlotte123):

LOL

OpenStudy (happinessbreaksbones):

LOL

OpenStudy (anonymous):

thanks for helping!

OpenStudy (shamil98):

\[\huge \frak You~are~very~welcome. ~~=)\]